The height of a box is 5 inches. The length is three inches more than the width. Find the width if the volume is 350.
Answer by shweta(56)

The box is a cuboid
height,h=5 inches, width=w and length,l=w+3 inches
Volume = 350
Volume =l*w*h
=(w+3)*w*5
=5w(w+3)
=5(w^2+ 3w)
350 = 5w^2 +15w
5w^2 +15w -350=0
Divide the equation by the H.C.F,5
w^2 + 3w -70=0
factors of 70= 2,5,7
w^2 + 10w -7w -70=0
w(w+10) -7(w+10)=0
(w+10)(w-7)=0
w=7,-10
the width of the cuboid cannot be negative ,so width is 7 inches and
length=w+3
l=7+3=10 inches